A movie theater charges $9 for adults and $7 for senior citizens. On a day when 325 people paid admission, the total receipts were $2495. How many people who paid were adults? How many were senior citizens?

Answer by mananth(16949)

Senior citizens tickets x Numbers
Adult ticket s ------------ y Numbers
x+y=325 .............1
7x+9y=2495.............2
Eliminate y
multiply (1)by -9
Multiply (2) by 1
-9x-9y=-2925
7x+9y=2495
Add the two equations
-2x = -430
/-2
x=215 Senior citizens tickets